Circular No. 7386

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

COMETS C/1999 J6, C/1999 U5, C/1999 W1, C/1999 Y3 Further to IAUC 7376 and 7383, D. A. Biesecker reports observations

(measures by D. Hammer and himself, reductions by B. G. Marsden) of four more comets seen in 1999 SOHO data, all but C/1999 J6 being presumed Kreutz sungrazers. Comets C/1999 J6 (visible in both C2 and C3 data) and C/1999 U5 (visible in only the C3 telescope) show no tail. However, C/1999 W1 and C/1999 Y3, which were both visible with only the C2 telescope, did show tails. Selected V magnitudes from Biesecker for C/1999 J6: May 10.750 UT, 8.1; 10.833, 7.3; 10.935, 6.5; 11.088, 5.9; 11.269, 5.5; 11.338, 4.9; 11.462, 5.1. Comet C/1999 J6 was first noted by M. Oates via the SOHO Web page; C/1999 U5 and C/1999 W1 were discovered by Biesecker, and C/1999 Y3 was discovered by A. Vourlidas of the SOHO team. Full astrometry and orbital elements appear on the MPECs indicated below.

1999 UT  α_{2000}  δ_{2000}  MPEC  Comet
May 10.726  3^{h}05.^{m}0 +15^{o}48′ 2000-F30  C/1999 J6
Oct. 31.495 14 06.8 −17 27  2000-F31  C/1999 U5
Nov. 23.188 15 55.1 −22 06  2000-F32  C/1999 W1
Dec. 21.496 18 04.0 −25 03  2000-F33  C/1999 Y3

COMETS C/2000 B5, C/2000 B6, C/2000 B7, C/2000 D3 Biesecker also reports observations, measured and reduced as above,

of four additional Kreutz sungrazing comets detected by SOHO in the past couple of months. Comet C/2000 B5 was discovered by Biesecker; the other three comets were found by several people browsing the SOHO web site, as follows: C/2000 B6, M. Meyer, T. Lovejoy, J. Shanklin, and M. Oates; C/2000 B7, Meyer and Lovejoy; C/2000 D3, Meyer and K. Černis. C/2000 D3 was visible with both the C2 and C3 telescopes and showed a short tail; the other three comets showed no apparent tail and were visible only in the C3 data. As before, full astrometry and orbital elements appear on the MPECs indicated below.

2000 UT  α_{2000}  δ_{2000}  MPEC  Comet
Jan. 17.863 20^{h}17.^{m}3 −23^{o}30′ 2000-F34  C/2000 B5
29.446 21 05.1 −19 55  2000-F35  C/2000 B6
31.846 21 12.4 −18 52  2000-F36  C/2000 B7
Feb. 26.413 22 42.2 − 9 34  2000-F37  C/2000 D3
2000 March 24  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Daniel W. E. Green